Solve by completing the square.
4x^2+2x-3=0
Thank you.

Answer by mukhopadhyay(490)

4x^2+2x-3 = 0
=> [(2x)^2 + 2*2x*(1/2) + (1/2)^2 - (1/2)^2] - 3 = 0
=> (2x+1/2)^2 = 3+1/4
=> (2x+1/2)^2 = 13/4
=> (2x+1/2) = sqrt(13)/2 or (2x+1/2) = -sqrt(13)/2
=> x = -1/4 + sqrt(13)/4 or x = -1/4 - sqrt(13)/4